OpenStudy (anonymous):

Finite Field / Over Zp Z_{12}=[0,1,2...,11] 7+6=1 3*5=3 7*8=8 5+2=? 6+5=?

OpenStudy (gorv):

what is this ???

OpenStudy (anonymous):

Over Zp systems.

OpenStudy (anonymous):

\[Z_3 ={0,1,2}\] 2+1=0 -1=2